Consider an old-fashioned bicycle with a small whe
Answer #1

we have to calculate this problem, first distance covered. _To_ solve the total I distance covered in constant acceleration fIL _distance covered during _deceleration. us_initial velocity - = 15.6 m/s ve final velocity =0 m/s t = time taken = 4.1 minCircumference of small wheel = 278 = 271 X 0.12 __m = 0.7536 m No. of revolutions made by small wheel Total distance coveredIf you have doubt in any part of the solution, feel free to ask.
Answer #2

SOLUTION :


v = 15.6 m/s in 1 min (60sec) 

Acceleration, a = ∆v / t = (15-6 - 0) / 60 = 0.26 m/s^2


Distance covered during constant acceleration , s1 

= average velocity * time

= (15.6 - 0)/2 * 60

= 468 m 


Distance covered at constant speed for 10.9 min, s2

= speed * time

= 15.6 * 10.9 * 60

= 10202.40 m


Time of deceleration = 4.1 min (246 sec)


Average speed during deceleration = 15.6 / 2 = 7.8 m/s


Distance covered during deceleration, s3

= average speed * time

= 7.8 * 246

= 1918.80 m 


Total distance covered, s

= s1 + s2 + s3

= 468 + 10202.40 + 1918.80 

= 12589.20 m 


Number of revolutions made by small wheel, N = s / (pi d) 


Small wheel turns angle through for this 

= N * 2 pi 

= s/(pi d) * 2pi 

= 2s/d

= 2*12589.20 / (0.12*2)

= 104910  rad. (ANSWER)

answered by: Tulsiram Garg